# Conditional Wasserstein Distances with Applications in Bayesian OT Flow Matching

### v2

@@ 8.3 Class conditional image generation
We apply our Bayesian OT flow matching for conditional image generation. We choose the condition $Y$ to be the class labels in order to generate samples of CIFAR10 for a given class. We simulate the flows for different values of $\beta$, by which we mean that we rescale the $Y$-component by $\beta$ as mentioned in Remark 11. We also simulate flows using the "diagonal" plans which coincide with the diagonal Bayesian flow matching objective [58]. For inference we simulate the flow ODE [17] using an adaptive step size solver (Runge-Kutta of order 5) . The samples in Fig. 3 are generated using the adaptive step size solver and sorted by class labels. For low values of $\beta$ we see that the resulting samples do not match their class labels, increasing $\beta$ leads to accurate class representations. The samples are generated given the labels of the training samples, therefore we see improved FID results as $\beta$ increases. The diagonal flow matching objective has the correct class representations, however since the associated couplings are not optimal our experiments suggest that this leads to higher variance during training and therefore slightly lower image quality, see [55] for more details on the advantages of OT based flow matching. We run each method for 500 epochs and compute an FID over the validation set every 20 epochs. Then for each method we choose the best checkpoint and report the results. The code is written in PyTorch [47] and is available online [^1] https://github.com/JChemseddine/Conditional_Wasserstein_Distances.
@@ 9 Conclusions
Inspired from applications in Bayesian inverse problems, we introduced conditional Wasserstein distances. We managed to rewrite these distances as expectations of the Wasserstein distances with respect to the observation. Therefore we are able to directly infer posterior guarantees in expectation when trained with the corresponding losses. Furthermore, we calculated the dual of the conditional Wasserstein-1 distance, when the probability measures are compactly supported and recovered well-known conditional Wasserstein GAN losses. We established corresponding velocity fields in the gradient flow theory for geodesics and used our results to design a new Bayesian flow matching algorithm. Moreover we use an approximation for the conditional Wasserstein distance depending on a parameter $\beta$ and show numerically that, when using it for class conditional flow matching, the result does respect the classes for sufficiently large $\beta$. We achieve better FID than random Bayesian flow matching on Cifar10. Future work includes improving Proposition 8 for non-discrete $P_{Y}$ distributions, where the measurability of the "conditional Monge maps" is the main obstruction. Further, from a practical viewpoint, a clear use case would be in conditional domain translation, i.e., when the latent distribution is not a standard Gaussian, but given by some data distribution. There, finding a good OT matching and making use of our proposed framework could improve existing algorithms.
Acknowledgement: P. Hagemann acknowledges funding from from the DFG within the project SPP 2298 "Theoretical Foundations of Deep Learning", C. Wald and G. Steidl gratefully acknowledge funding by the DFG within the SFB “Tomography Across the Scales” (STE 571/19-1, project number: 495365311).
@@ Appendix A Counterexample for equality in equation 2
where $\mathcal{S}_{n}$ is the set of permutations on $\{1,\ldots,n\}$, see [48, Proposition 2.1].
@@
Note that in [35], the summation metric is considered, i.e. $\|(x_{1},y_{1})-(x_{2},y_{2})\|_{sum}=\|x_{1}-x_{2}\|+\|y_{1}-y_{2}\|$ for which our counterexample is still valid.
@@ Appendix B Proofs of Section 3
Proof of Proposition 1. iii) Let $\alpha$ be defined by (11) which was already used in the proof of [35, Theorem 2] , i.e.,
@@
for all Borel measurable functions $f:(A\times B)^{2}\to[0,+\infty]$. Indeed $\alpha$ is a well defined probability measure on $(A\times B)^{2}$ by the following reasons: by [6, Lemma 12.4.7], we can choose a Borel family $(\alpha_{y})_{y}$. For any Borel set $\mathcal{S}\subseteq A\times B\times B$, we have
@@
By [6, Equation 5.3.1] the function $y\mapsto\int_{B^{2}}1_{\mathcal{S}}(y,x_{1},x_{2})\mathrm{d}\alpha_{y}$ is Borel measurable. Consequently also $y\mapsto\delta_{y}\times\alpha_{y}(\mathcal{S})$ is Borel measurable and thus $\alpha$ is well defined.
@@
$=\int_{A}\int_{B^{2}}\|x_{1}-x_{2}\|^{p}\,\mathrm{d}\pi^{2,3,4}_{\sharp}\alpha.\qquad\Box$
@@ Appendix C Proofs of Section 4
The proof uses similar arguments as the short notes [54] and [11], which are derivations for the dual for the "usual" Wasserstein distance. We adapt these ideas for our conditional Wasserstein distance.
Proof of Proposition 4. Let $C_{b}=C_{b}(A\times B)$ be the space of continuous bounded functions on $A\times B$ and $\mathcal{S}$ the set of nonnegative finite Borel measures $\alpha$ on $(A\times B)^{2}$ which are supported at most on the y-diagonal. By [52, Section 1.2], we know that
@@
for all $y\in A$, since otherwise the attained infimum is $\infty$- ∞. Therefore we have for the optimal $f,g$ that $L(\alpha,f,g)\geq\mathbb{E}_{Y,X}[f]+\mathbb{E}_{Y,Z}[g]$ and choosing the plan $\alpha=0\in\mathcal{S}$, we obtain
@@
~:=$\tilde{\mathcal{F}}:=\{(f,g)\in(C_{b}(A\times B))^{2}:f(y,x_{1})+g(y,x_{2})\leq\|x_{2}-x_{2}\|\}.$
@@
Thus, $W_{1,Y}(P_{Y,X},P_{Y,Z})=\sup_{h\in\mathcal{F}}\{\mathbb{E}_{Y,X}[h]-\mathbb{E}_{Y,Z}[h]\}$, which finishes the proof. $\Box$
The proof of strong duality relies on the following minimax principle from [9, Theorem 7 Chapter 6].
Theorem 14.
@@
Proof.
We will verify the conditions in Theorem 14. Recall that $\mathcal{S}$ is the set of finite nonnegative Borel measures $\alpha$ on $(A\times B)^{2}$ such that there exists a finite nonegative finite measure $\beta$ on $B$ with $\pi^{1,3}_{\sharp}\alpha=\Delta_{\sharp}\beta$. Let $\mathcal{M}$ be the topological vector space of finite signed Borel measures on $(A\times B)^{2}$ with weak convergence topology. Thus, since the pushforward is linear on $\mathcal{S}$, we conclude that $\mathcal{S}$ is a convex subset. Now we use Theorem 14 with $X\coloneqq\mathcal{S}$, $Y\coloneqq C_{b}\times C_{b}$ and $f\coloneqq L$.
@@
can become arbitrary large which contradicts $L(\alpha,-1,-1)\leq a$. Therefore the compactness of $A,B$ implies that $\mathcal{S}_{a}$ is a family of tight measures. By [12, Theorem 8.6.7], the set $\mathcal{S}_{a}$ is relatively compact in the weak topology. ∎
@@ Appendix D Proofs of Section 5
Finally, the desired equality follows like in [6, Theorem 7.2.2]for $0\leq s\leq t\leq 1$ by
@@
iii) Recall (see [6, Section 5.1]) that a sequence $\mu_{k}\in\mathcal{P}(\mathbb{R}^{n})$ is said to converge weakly to $\mu\in\mathcal{P}(\mathbb{R}^{n})$ if $\lim_{k\to\infty}\int_{\mathbb{R}^{n}}f(x)\,\mathrm{d}\mu_{k}(x)=\int_{\mathbb{R}^{n}}f(x)\,\mathrm{d}\mu(x)$ for all $f\in C_{b}(\mathbb{R}^{n})$. By the dominated convergence theorem, we have for $\mu_{s}=(e_{s})_{\sharp}\alpha$ and every $f\in C_{b}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ that
@@
which finishes the proof. $\Box$
Proof of Proposition 6. The proof of $i),ii),iv)$ is almost identical to the proofs of [5, Theorem 17.2, Lemma 17.3]. For the convenience of the reader and because we also need measurability of $v_{t}$ in $t$ we include a slight adaptation of their proof. We let $e:[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}\times\mathbb{R}^{d}\times\mathbb{R}^{m}\to[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}$ be defined by $e(t,y_{1},x_{1},y_{2},x_{2})=(t,(1-t)(y_{1},x_{1})+t(y_{2},x_{2}))$. For $\mathcal{L}$ the Lebesgue measure we have that
@@
for every bounded $\mathcal{L}\otimes\alpha$ measurable function $f$. In particular using $f=1_{B}$ for a Borel set $B\in\mathcal{B}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ we see that $\mu_{t}:[0,1]\times\mathcal{B}\to\mathbb{R}$ is a Markov kernel, i.e. the measure $\int_{0}^{1}\mu_{t}\mathrm{d}t$ is well defined, and $e_{\sharp}\left(\mathcal{L}\otimes\alpha\right)=\int_{0}^{1}\mu_{t}\mathrm{d}t$. Hence [5, Lemma 17.3] (with $e\coloneqq e$, $\mu=\mathcal{L}\otimes\alpha$, $v=(y_{2},x_{2})-(y_{1},x_{1})$, $w=v$) implies that there exists $v\in L^{2}_{\int\mu_{t}\mathrm{d}t}([0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m})$ such that
@@
we obtain from (77) that $e_{t,\sharp}\left(((y_{2},x_{2})-(y_{1},x_{1}))\alpha\right)=v(t,y,x)\mu_{t}$ for $\mathcal{L}$ a.e. $t\in[0,1]$ which shows $i)$. By [5, Lemma 17.3] we obtain that
@@
since $\varphi(1,\cdot)=\varphi(0,\cdot)=0$ because $\varphi$ is compactly supported on $(0,1)\times\mathbb{R}^{d}\times\mathbb{R}^{m}$. $\Box$
@@
    $\mu_{0},\mu_{1}$ both admit densities $\rho_{0},\rho_{1}$ which are supported on open, convex, bounded, connected subsets $\Omega_{0},\Omega_{1}\subset\mathbb{R}^{m}$ on which they are bounded away from 00 and $\infty$. Assume further that $\rho_{0}\in C^{2}(\Omega_{0}),\rho_{1}\in C^{2}(\Omega_{1})$. Let $T$ be the optimal Monge map with associated optimal plan $\alpha\in\Gamma(\mu_{0},\mu_{1})$.
@@
Proof.
i): Let $\mu_{0}=\frac{1}{n}\sum_{i=1}^{n}\delta_{a_{i}},\mu_{1}=\frac{1}{n}\sum_{i=1}^{n}\delta_{b_{i}}$ and let $T$ be a optimal map. The associated optimal plan is then $\alpha=\frac{1}{n}\sum_{i=1}^{n}\delta_{a_{i},T(a_{i})}$. Using $e_{t,\sharp}(x_{2}-x_{1})\alpha=v_{t}\mu_{t}$ and $\mu_{t}=\frac{1}{n}\sum_{i=1}^{n}\delta_{T_{t}(a_{i})}$ for $T_{t}(x)=(1-t)x+tT(x)$ we can conclude
@@
ii): First, note that by [5, (16.12)] if there exists an invertible Monge map $T$ then the geodesic between $\mu_{0},\mu_{1}$ fulfills the continuity equation with vector field
@@
where $T_{t}=(1-t)\text{Id}+tT$. By Caffarelli’s regularity Theorem [56, Theorem 12.5, ii)], we get the existence of a unique Monge map $T\in C^{1}(\Omega_{0})$ mapping $\mu_{0}$ to $\mu_{1}$ and $U\in C^{1}(\Omega_{1})$ mapping $\mu_{1}$ to $\mu_{0}$. By [5, Theorem 5.2] we know that $T\circ U=\text{Id}$ on $\Omega_{1}$ and $U\circ T=\text{Id}$ on $\Omega_{0}$ and thus $T:\Omega_{0}\to\Omega_{1}$ is a $C^{1}$ diffeomorphism and in particular $\det(\nabla T)\neq 0$ on $\Omega_{0}$. Since we know by [6, Proposition 6.2.12] that $\nabla T$ is positive definite $\mu_{1}$ a.e. on $\Omega_{0}$ we can deduce from $\det(\nabla T)\neq 0$ that $\nabla T$ is positive definite on $\Omega_{0}$. Consequently for $T_{t}=(1-t)\text{Id}+tT$ we have that $\nabla T_{t}=(1-t)\text{Id}+t\nabla T$ is positive definite on $\Omega_{0}$ and thus the image of $\Omega_{0}$ under $T_{t}$ is open. Furthermore, we know by the proof of [6, Proposition 6.2.12] that $T_{t}$ as a Monge map from $\mu_{0}$ to $\mu_{t}$ is injective on all points where $\nabla T_{t}$ is positive definite, which is on the whole $\Omega_{0}$, and thus $T_{t}$ is a diffeomorphism onto its image. Consequently, it possesses a $C^{1}$ inverse $T_{t}^{-1}:T_{t}(\Omega_{0})\to\Omega_{0}$. Furthermore $(t,x)\mapsto(t,T_{t}(x))$ is an bijective Borel map from $[0,1]\times\Omega_{0}$ onto its Borel measurable image which we denote by $\Omega\subset[0,1]\times\mathbb{R}^{m}$. Thus $T_{t}^{-1}$ is a Borel measurable map from $\Omega\to\Omega_{0}$. Then for $v_{t}:=(T-\text{Id})\circ T_{t}^{-1}:T_{t}(\Omega_{0})\to\mathbb{R}^{m}$ we have that $v_{t}$ is Borel measurable on $\Omega$ and also $\phi_{t}=T_{t}:\Omega_{0}\to\mathbb{R}^{m}$ is Borel measurable. Furthermore, we have
@@
Since we can set $\phi_{t}(x)=x$ on $\mathbb{R}^{m}\setminus\Omega_{0}$ and $v_{t}(x)=0$ for $x\in\mathbb{R}^{m}\setminus T_{t}(\Omega_{0})$, we obtain the claim. ∎
@@
for $\mu_{0}$-a.e. $(y,x)\in\mathbb{R}^{d}\times\mathbb{R}^{m}$. Note that since $\tilde{v}_{t,y_{i}}$ is Borel measurable on $[0,1]\times\mathbb{R}^{m}$ we can assume that $\tilde{v}_{t}$ is Borel measurable on $[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}$ and similarly for $\phi_{t}$. Furthermore

### v3

@@ 8.3 Class Conditional Image Generation
We apply our Bayesian OT flow matching for conditional image generation. We choose the condition $Y$ to be the class labels in order to generate samples of CIFAR10 for a given class. We simulate the flows for different values of $\beta$, by which we mean that we rescale the $Y$-component by $\beta$ as mentioned in Remark 11. We also simulate flows using the "diagonal" plans which coincide with the diagonal Bayesian flow matching objective Wildberger et al. (2023). For inference we simulate the flow ODE Chen (2018) using an adaptive step size solver (Runge-Kutta of order 5) . The samples in Fig. 3 are generated using the adaptive step size solver and sorted by class labels. For low values of $\beta$ we see that the resulting samples do not match their class labels, increasing $\beta$ leads to accurate class representations. The samples are generated given the labels of the training samples, therefore we see improved FID results as $\beta$ increases. The diagonal flow matching objective has the correct class representations, however since the associated couplings are not optimal our experiments suggest that this leads to higher variance during training and therefore slightly lower image quality, see Tong et al. (2023) for more details on the advantages of OT based flow matching. We run each method for 500 epochs and compute an FID over the validation set every 20 epochs. Then for each method we choose the best checkpoint and report the results. The code is written in PyTorch Paszke et al. (2019) and is available online [^1] https://github.com/JChemseddine/Conditional_Wasserstein_Distances.
@@ 9 Conclusions
Inspired from applications in Bayesian inverse problems, we introduced conditional Wasserstein distances. We managed to rewrite these distances as expectations of the Wasserstein distances with respect to the observation. Therefore we are able to directly infer posterior guarantees in expectation when trained with the corresponding losses. Furthermore, we calculated the dual of the conditional Wasserstein-1 distance, when the probability measures are compactly supported and recovered well-known conditional Wasserstein GAN losses. We established corresponding velocity fields for geodesics and used our results to design a new Bayesian flow matching algorithm. Moreover we use an approximation for the conditional Wasserstein distance depending on a parameter $\beta$ and show numerically that, when using it for class conditional flow matching, the result does respect the classes for sufficiently large $\beta$. We achieve better FID than random Bayesian flow matching on Cifar10. Future work includes conditional domain translation, i.e., when the latent distribution is not a standard Gaussian, but given by some data distribution. There, finding a good OT matching and making use of our proposed framework could improve existing algorithms.
Acknowledgments and Disclosure of Funding
Many thanks to F. Vialard for pointing to reference Peszek and Poyato (2023) after reading our arXiv version. P. Hagemann acknowledges funding from from the DFG within the project SPP 2298 "Theoretical Foundations of Deep Learning", C. Wald and G. Steidl gratefully acknowledge funding by the DFG within the SFB “Tomography Across the Scales” (STE 571/19-1, project number: 495365311) and by the DFG within the Exellencecluster MATH+.
@@ Appendix A Counterexample for Equality in Equation 2
where $\mathcal{S}_{n}$ is the set of permutations on $\{1,\ldots,n\}$, see (Peyré and Cuturi, 2019, Proposition 2.1).
@@
Note that in (Kim et al., 2023), the summation metric is considered, i.e. $\|(x_{1},y_{1})-(x_{2},y_{2})\|_{sum}=\|x_{1}-x_{2}\|+\|y_{1}-y_{2}\|$ for which our counterexample is still valid.
@@ Appendix B Proofs of Section 3
Proof of Proposition 1. iii) Let $\alpha$ be defined by (11) which was already used in the proof of (Kim et al., 2023, Theorem 2) , i.e.,
@@
for all Borel measurable functions $f:(A\times B)^{2}\to[0,+\infty]$. Indeed $\alpha$ is a well defined probability measure on $(A\times B)^{2}$ by the following reasons: by (Ambrosio et al., 2005, Lemma 12.4.7), we can choose a Borel family $(\alpha_{y})_{y}$. For any Borel set $\mathcal{S}\subseteq A\times B\times B$, we have
@@
By (Ambrosio et al., 2005, Equation 5.3.1) the function $y\mapsto\int_{B^{2}}1_{\mathcal{S}}(y,x_{1},x_{2})\mathrm{d}\alpha_{y}$ is Borel measurable. Consequently also $y\mapsto\delta_{y}\times\alpha_{y}(\mathcal{S})$ is Borel measurable and thus $\alpha$ is well defined.
@@
$=\int_{A}\int_{B^{2}}\|x_{1}-x_{2}\|^{p}\,\mathrm{d}\pi^{2,3,4}_{\sharp}\alpha.\qquad\blacksquare$
@@ Appendix C Proofs of Section 4
The proof uses similar arguments as the short notes (Thickstun, ) and (Basso, 2015), which are derivations for the dual for the "usual" Wasserstein distance. We adapt these ideas for our conditional Wasserstein distance.
Proof of Proposition 4. Let $C_{b}=C_{b}(A\times B)$ be the space of continuous bounded functions on $A\times B$ and $\mathcal{S}$ the set of nonnegative finite Borel measures $\alpha$ on $(A\times B)^{2}$ which are supported at most on the y-diagonal. By (Santambrogio, 2015, Section 1.2), we know that
@@
for all $y\in A$, since otherwise the attained infimum is -$\infty$. Therefore we have for the optimal $f,g$ that $L(\alpha,f,g)\geq\mathbb{E}_{Y,X}[f]+\mathbb{E}_{Y,Z}[g]$ and choosing the plan $\alpha=0\in\mathcal{S}$, we obtain
@@
$\tilde{\mathcal{F}}:=\{(f,g)\in(C_{b}(A\times B))^{2}:f(y,x_{1})+g(y,x_{2})\leq\|x_{2}-x_{2}\|\}.$
@@
Thus, $W_{1,Y}(P_{Y,X},P_{Y,Z})=\sup_{h\in\mathcal{F}}\{\mathbb{E}_{Y,X}[h]-\mathbb{E}_{Y,Z}[h]\}$, which finishes the proof. $\blacksquare$
The proof of strong duality relies on the following minimax principle from (Aubin and Ekeland, 2006, Theorem 7 Chapter 6).
Theorem 14
@@
Proof  We will verify the conditions in Theorem 14. Recall that $\mathcal{S}$ is the set of finite nonnegative Borel measures $\alpha$ on $(A\times B)^{2}$ such that there exists a finite nonegative finite measure $\beta$ on $B$ with $\pi^{1,3}_{\sharp}\alpha=\Delta_{\sharp}\beta$. Let $\mathcal{M}$ be the topological vector space of finite signed Borel measures on $(A\times B)^{2}$ with weak convergence topology. Thus, since the pushforward is linear on $\mathcal{S}$, we conclude that $\mathcal{S}$ is a convex subset. Now we use Theorem 14 with $X\coloneqq\mathcal{S}$, $Y\coloneqq C_{b}\times C_{b}$ and $f\coloneqq L$.
@@
can become arbitrary large which contradicts $L(\alpha,-1,-1)\leq a$. Therefore the compactness of $A,B$ implies that $\mathcal{S}_{a}$ is a family of tight measures. By (Bogachev, 2007, Theorem 8.6.7), the set $\mathcal{S}_{a}$ is relatively compact in the weak topology.
@@ Appendix D Proofs of Section 5
Finally, the desired equality follows like in (Ambrosio et al., 2005, Theorem 7.2.2)for $0\leq s\leq t\leq 1$ by
@@
iii) Recall (see (Ambrosio et al., 2005, Section 5.1)) that a sequence $\mu_{k}\in\mathcal{P}(\mathbb{R}^{n})$ is said to converge weakly to $\mu\in\mathcal{P}(\mathbb{R}^{n})$ if $\lim_{k\to\infty}\int_{\mathbb{R}^{n}}f(x)\,\mathrm{d}\mu_{k}(x)=\int_{\mathbb{R}^{n}}f(x)\,\mathrm{d}\mu(x)$ for all $f\in C_{b}(\mathbb{R}^{n})$. By the dominated convergence theorem, we have for $\mu_{s}=(e_{s})_{\sharp}\alpha$ and every $f\in C_{b}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ that
@@
which finishes the proof. $\blacksquare$
Proof of Proposition 6. The proof of $i),ii),iv)$ is almost identical to the proofs of (Ambrosio et al., 2021, Theorem 17.2, Lemma 17.3). For the convenience of the reader and because we also need measurability of $v_{t}$ in $t$ we include a slight adaptation of their proof. We let $e:[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}\times\mathbb{R}^{d}\times\mathbb{R}^{m}\to[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}$ be defined by $e(t,y_{1},x_{1},y_{2},x_{2})=(t,(1-t)(y_{1},x_{1})+t(y_{2},x_{2}))$. For $\mathcal{L}$ the Lebesgue measure we have that
@@
for every bounded $\mathcal{L}\otimes\alpha$ measurable function $f$. In particular using $f=1_{B}$ for a Borel set $B\in\mathcal{B}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ we see that $\mu_{t}:[0,1]\times\mathcal{B}\to\mathbb{R}$ is a Markov kernel, i.e. the measure $\int_{0}^{1}\mu_{t}\mathrm{d}t$ is well defined, and $e_{\sharp}\left(\mathcal{L}\otimes\alpha\right)=\int_{0}^{1}\mu_{t}\mathrm{d}t$. Hence (Ambrosio et al., 2021, Lemma 17.3) (with $e\coloneqq e$, $\mu=\mathcal{L}\otimes\alpha$, $v=(y_{2},x_{2})-(y_{1},x_{1})$, $w=v$) implies that there exists $v\in L^{2}_{\int\mu_{t}\mathrm{d}t}([0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m})$ such that
@@
we obtain from (77) that $e_{t,\sharp}\left(((y_{2},x_{2})-(y_{1},x_{1}))\alpha\right)=v(t,y,x)\mu_{t}$ for $\mathcal{L}$ a.e. $t\in[0,1]$ which shows $i)$. By (Ambrosio et al., 2021, Lemma 17.3) we obtain that
@@
since $\varphi(1,\cdot)=\varphi(0,\cdot)=0$ because $\varphi$ is compactly supported on $(0,1)\times\mathbb{R}^{d}\times\mathbb{R}^{m}$. $\blacksquare$
@@
    $\mu_{0},\mu_{1}$ both admit densities $\rho_{0},\rho_{1}$ which are supported on open, convex, bounded, connected subsets $\Omega_{0},\Omega_{1}\subset\mathbb{R}^{m}$ on which they are bounded away from 0 and $\infty$. Assume further that $\rho_{0}\in C^{2}(\Omega_{0}),\rho_{1}\in C^{2}(\Omega_{1})$. Let $T$ be the optimal Monge map with associated optimal plan $\alpha\in\Gamma(\mu_{0},\mu_{1})$.
@@
Proof  i): Let $\mu_{0}=\frac{1}{n}\sum_{i=1}^{n}\delta_{a_{i}},\mu_{1}=\frac{1}{n}\sum_{i=1}^{n}\delta_{b_{i}}$ and let $T$ be a optimal map. The associated optimal plan is then $\alpha=\frac{1}{n}\sum_{i=1}^{n}\delta_{a_{i},T(a_{i})}$. Using $e_{t,\sharp}(x_{2}-x_{1})\alpha=v_{t}\mu_{t}$ and $\mu_{t}=\frac{1}{n}\sum_{i=1}^{n}\delta_{T_{t}(a_{i})}$ for $T_{t}(x)=(1-t)x+tT(x)$ we can conclude
@@
ii): First, note that by (Ambrosio et al., 2021, (16.12)) if there exists an invertible Monge map $T$ then the geodesic between $\mu_{0},\mu_{1}$ fulfills the continuity equation with vector field
@@
where $T_{t}=(1-t)\text{Id}+tT$. By Caffarelli’s regularity Theorem (Villani, 2009, Theorem 12.50, ii)), we get the existence of a unique Monge map $T\in C^{1}(\Omega_{0})$ mapping $\mu_{0}$ to $\mu_{1}$ and $U\in C^{1}(\Omega_{1})$ mapping $\mu_{1}$ to $\mu_{0}$. By (Ambrosio et al., 2021, Theorem 5.2) we know that $T\circ U=\text{Id}$ on $\Omega_{1}$ and $U\circ T=\text{Id}$ on $\Omega_{0}$ and thus $T:\Omega_{0}\to\Omega_{1}$ is a $C^{1}$ diffeomorphism and in particular $\det(\nabla T)\neq 0$ on $\Omega_{0}$. Since we know by (Ambrosio et al., 2005, Proposition 6.2.12) that $\nabla T$ is positive definite $\mu_{1}$ a.e. on $\Omega_{0}$ we can deduce from $\det(\nabla T)\neq 0$ that $\nabla T$ is positive definite on $\Omega_{0}$. Consequently for $T_{t}=(1-t)\text{Id}+tT$ we have that $\nabla T_{t}=(1-t)\text{Id}+t\nabla T$ is positive definite on $\Omega_{0}$ and thus the image of $\Omega_{0}$ under $T_{t}$ is open. Furthermore, we know by the proof of (Ambrosio et al., 2005, Proposition 6.2.12) that $T_{t}$ as a Monge map from $\mu_{0}$ to $\mu_{t}$ is injective on all points where $\nabla T_{t}$ is positive definite, which is on the whole $\Omega_{0}$, and thus $T_{t}$ is a diffeomorphism onto its image. Consequently, it possesses a $C^{1}$ inverse $T_{t}^{-1}:T_{t}(\Omega_{0})\to\Omega_{0}$. Furthermore $(t,x)\mapsto(t,T_{t}(x))$ is an bijective Borel map from $0,1]\times\Omega_{0}$ onto its Borel measurable image which we denote by $\Omega\subset[0,1]\times\mathbb{R}^{m}$. Thus $T_{t}^{-1}$ is a Borel measurable map from $\Omega\to\Omega_{0}$. Then for $v_{t}:=(T-\text{Id})\circ T_{t}^{-1}:T_{t}(\Omega_{0})\to\mathbb{R}^{m}$ we have that $v_{t}$ is Borel measurable on $\Omega$ and also $\phi_{t}=T_{t}:\Omega_{0}\to\mathbb{R}^{m}$ is Borel measurable. Furthermore, we have
@@
Since we can set $\phi_{t}(x)=x$ on $\mathbb{R}^{m}\setminus\Omega_{0}$ and $v_{t}(x)=0$ for $x\in\mathbb{R}^{m}\setminus T_{t}(\Omega_{0})$, we obtain the claim.
@@
for $\mu_{0}$-a.e. $(y,x)\in\mathbb{R}^{d}\times\mathbb{R}^{m}$. Note that since $\tilde{v}_{t,y_{i}}$ is Borel measurable on $0,1]\times\mathbb{R}^{m}$ we can assume that $\tilde{v}_{t}$ is Borel measurable on $0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}$ and similarly for $\phi_{t}$. Furthermore